# Forgetting Order of Continual Learning: Examples That are Learned First are Forgotten Last

### v1

@@ 2.4 How the replay buffer changes the catastrophic forgetting dynamics
Figure 3: Impact of the buffer size on forgetting dynamics. (a-c) Similarly to Fig. 2(c), the first task’s binary matrices $M$, when adding replay buffer of varying buffer sizes ($4\%$, $12\%$, and $40\%$). Rows denote examples, and columns denote epochs, indicating correct classifications by the majority of $10$ networks. Rows are sorted by learning speed. In all cases, examples that were learned faster are more likely to be remembered. With bigger buffers, the networks can remember gradually slower-to-learn examples. (d) The mean learning speed of remembered examples of models with different buffer sizes. Models with bigger replay buffers remember slower-to-learn examples.
To evaluate replay-based methods, we incorporate a replay buffer in our experiments. At the end of the first task, we uniformly sample examples to a fixed-memory replay buffer. During training on the subsequent task, we adopt the experience replay technique [46], alternating batch sampling between the new data and the contents of the replay buffer.
@@
Similarly to §2, we trained $10$ networks on CIFAR-100-2, with a replay buffers of varying sizes: $1000$, $3000$, and $10000$ samples corresponding to $4\%$, $12\%$, and $40\%$ of the data size, respectively. In Figs. 3(a-c), we visualize the aggregated epoch-wise classification matrix of the test data for each case. The rows in each matrix are sorted by the learning speed of the examples. As expected, integrating a replay buffer mitigates some catastrophic forgetting, boosting the model performance on the first task, with larger buffers yielding better results. Nevertheless, similar to the scenario without a replay buffer $B=\emptyset$, a relationship exists between learning speed and catastrophic forgetting: networks tend to remember fast-learned examples while forgetting those learned later even when using a uniformly sampled replay buffer. Notably, models with larger buffer sizes remember most of the examples that models with smaller buffer sizes do, while additionally remembering gradually slower-learned examples. Similar results occur when changing the network architecture (see App.B)
@@
We quantify the relationship between the buffer size and remembered examples’ learning speed. Training $10$ networks on CIFAR-100-2 with various buffer sizes, we plot the mean learning speed of examples remembered by the models for each buffer size. A robust correlation emerges ($r=0.966$, $p\leq 10^{-6}$): models with smaller buffer sizes remember quickly learned examples, whereas larger models with larger buffers enable the model also to remember slower-learned. These results are plotted in Fig. 3(d). Similar results were achieved across diverse datasets, architectures, hyperparameters, and continual learning algorithms, see App. B.
3 Effects of different replay buffer compositions on catastrophic forgetting
Replay-based continual learning methods typically sample the replay buffer uniformly, including both quickly and slowly learned examples. In this section, we analyze various replay buffer compositions and find that medium-speed learned examples are the most beneficial for the replay process. These examples, which the model has only a moderate chance of remembering, prove particularly important. One possible explanation is that increasing the buffer size impacts the catastrophic forgetting of these examples the most, suggesting that focusing on them is advantageous.
3.1 The Goldilocks approach – methodology and definitions
To explore different buffer compositions, we remove a combination of the quickest and slowest learned examples from the data and then sample the buffer uniformly from the remaining examples. This allows us to add a bias to the replay buffer toward examples learned at certain speeds. We propose Goldilocks, a novel sampling method, used both in our empirical analysis and as a practical replay buffer sampling technique. The name Goldilocks reflects the method’s focus on examples learned at an optimal speed, akin to the story where Goldilocks prefers things that are "just right."
Formally, Goldilocks is defined by two hyperparameters, $quick$ and $slow$, which specify the percentage of examples learned too quickly or too slowly to be removed from dataset $\mathcal{D}_{t}$. During training of the model $f$ on $\mathcal{D}_{t}$, we maintain the epoch-wise classification matrix $M$ (see § 2.2). At the end of task $t$, examples in $\mathcal{D}_{t}$ are sorted by learning speed, and the top %$\%quick$ and bottom %$\%slow$ are removed. The replay buffer is then uniformly sampled from the remaining examples. Notably, for tasks beyond the initial two, when reducing the replay buffer size, examples can be uniformly sampled out. The pseudocode for Goldilocks sampling is provided in Algorithm 1.
Algorithm 1 Training continual learning method $CL$ with Goldilocks sampling.
Input: task $\mathcal{D}_{t}$, sample size $|B|$, epochs $E$, amount of quick/slow learned examples to remove $q,s$. Output: buffer of size $|B|$
1:$Classification\_Matrix$leftarrow ${0}^{E\times|\mathcal{D}_{t}|}$
2:for $e=1,...,E$ do
3:     Train the model $f$ one epoch using the selected continual learning method $CL$
4:     for $(x_{i},y_{i})\in\mathcal{D}_{t}$ do $\triangleright$ Can be calculated efficiently in the forward pass
5:         Classification_Matrix$[e,i]\leftarrow\mathds{1}[f(x_{i})=y_{i}]$
6:     end for
7:end for
8:$Learning\_Speed$leftarrow Mean($Classification\_Matrix$, axis=0)
9:$\mathcal{D}_{t}^{{}^{\prime}}\leftarrow$ remove $q\%$ highest and $s\%$ lowest $Learning\_Speed$ examples from $\mathcal{D}_{t}$
10:return $|B|$ examples sampled uniformly from $\mathcal{D}_{t}^{{}^{\prime}}$
3.2 Which examples should appear in the replay buffer
(a) CIFAR-100-2, $4\%$ buffer
(b) CIFAR-100-2, $40\%$ buffer
(c) TinyImagenet-2, $20\%$ buffer
Figure 4: Comparing different replay buffer compositions. Each matrix compares the mean final test accuracy of 10 networks trained with different replay buffer compositions achieved by modifying Goldilocks’ $quick$ and $slow$ hyperparameters. The ($quick$, $slow$) = ($0\%$, $0\%$) case, marked by a black box, corresponds to uniform sampling. (a) CIFAR-100-2 with a $4\%$ buffer. Removing faster-to-learn examples is more beneficial. (b) CIFAR-100-2 with a $40\%$ buffer. Removing slower-to-learn examples is more beneficial. (c) TinyImageNet-2 with a $40\%$ buffer. Across all scenarios, focusing on examples learned in the middle of the learning process is most beneficial. Smaller buffers exhibit greater benefits from prioritizing faster-to-learn examples.
(d) (A)–>(B)
(e) (A)–>(C)
Figure 5: Comparing replay buffer compositions when training the same original task, followed by different subsequent tasks. We consider 3 tasks, A, B, and C, training networks on task A followed by either B or C. All tasks are subsets of CIFAR-100, where Task A is classes 0-24, B is classes 25-49 and C is classes 50-74. (a) training on task A followed by B. (b) training on task A followed by C. The results format is similar to Fig. 3.2. Despite changes in the subsequent tasks, effective and ineffective buffer compositions for the original task A remain consistent.
(a) GEM, CIFAR-100-2, $1\%$ buffer
Figure 6: Comparing different replay buffer compositions for GEM with a buffer of $1\%$. The experimental setup replicates the one done in Fig. 3.2. Focusing on examples learned midway through the learning process is most beneficial. Due to the smaller buffer, removing more easier examples is better.
We investigate the effect of constructing the replay buffer with examples learned at different stages of the learning process. This is achieved by adjusting the $slow$ and $fast$ parameters in Goldilocks (see §3.1). Regardless of these parameters’ values, the replay buffer maintains a constant size but is sampled from different data distributions. We evaluate a wide range of $slow$ and $fast$ values by training 10 networks per parameter set on various datasets and buffer sizes. The mean final test accuracy across all tasks is plotted in a grid-like heatmap, in Fig. 3.2. To aid visualization, standard errors are plotted separately, see App. C. Notably, the point $(slow,fast)=(0,0)$ (marked by a black box) represents training with a uniformly sampled replay buffer, as is done in the original works.
In Figs. 3.2(a-b), we compare replay buffer compositions for networks trained on CIFAR-100-2. We consider a small buffer of size $1000$ ($4\%$ of data size) and a large buffer of size $10000$ ($40\%$ of data size). Blank entries denote infeasible parameters where too much data is removed, rendering a $40\%$ buffer impossible. In both cases, removing both slowly- and quickly-learned examples proves beneficial, with parameters mutually influencing each other. Focusing on examples learned midway through the learning process, where the network can either remember or forget, is most effective in reducing catastrophic forgetting. When the buffer is smaller, optimal results are achieved with a bias toward fast-learned examples, as models struggle to remember simpler examples in such scenarios.
@@ 3.2 Which examples should appear in the replay buffer
In Fig. 4(c), we replicate this experiment with TinyImagenet-2, with a buffer size of $10000$ ($20\%$ of the data). Similarly to the CIFAR-100-2 case, focusing on examples learned midway through learning yields the best performances. It’s worth noting that across all cases, the range of hyperparameters that enhance performance is broad and continuous, suggesting that improvements are not confined to specific hyperparameter sets; instead, various methods of focusing on midway-learned examples yield positive outcomes. Similar results with class incremental learning can be found in App. A.
3.3 Independency of the optimal buffer composition on subsequent tasks
@@
We initially focus on tasks from the same datasets. We pick task A to be the first 25 classes of CIFAR-100, task B to be classes 26 to 50, and task C to be classes 51 to 75. In Fig. 5, we depict the mean final accuracy of 10 networks across various replay buffer compositions when using either task B or task C as the subsequent task after task A. Despite some quantitative variations in performance, the qualitative results consistently indicate that good buffer compositions for one case remain effective for the other, and vice versa. These results suggest that the ideal replay buffer composition of task A is robust to differences between tasks B and C in this context.
@@
Further, we explore scenarios involving either changing the classification task or replacing it with memorization by training on random labels. We adopt the RotNet approach [13] to change the classification task. We pick tasks A and B to be different subsets of CIFAR-100, while task C involves a rotation classification task, where each example in task A is randomly rotated between $90\degree,180\degree$, or $270\degree$ degrees, with the task being to classify these rotations regardless of the original label. For the random case, tasks A and B remain unchanged, while task C involves the same subset of task B, but with labels chosen uniformly at random, as suggested in [57]. In both scenarios, we observe consistent results: the ideal replay buffer composition of task A remains robust to differences between tasks B and C in both contexts. Further details and heatmaps for the results are available in App. D.
Choosing hyperparameters. Implementing Goldilocks practically requires identifying the optimal $quick$ and $slow$ hyperparameters without extra data. While simple heuristics can often suffice due to the broad range of hyperparameters yielding good results, our findings propose an alternative approach for hyperparameter fine-tuning. By conducting a grid search for both parameters, training the network on the original task, and using the rotated version as the subsequent task, efficient parameter selection can be achieved. The smooth impact of parameters on the results implies that a coarse grid search is often adequate, reducing computational burden.
3.4 Goldilocks sampling improves various continual learning methods
Many replay-based continual learning algorithms sample the replay buffer uniformly. In this section, we investigate the impact of replacing uniform sampling with Goldilocks in various such algorithms, including DER [5], DER++ [5], X-DER [4], AGEM [8], ER [46], GEM [30], RPC [39], and FDR [51]. Our results demonstrate that Goldilocks consistently improves the performance of all tested continual algorithms, thereby advancing the state of the art in multiple image classification tasks.
We replicate the experimental setup outlined in §3.2, exploring various compositions of the replay buffer in a grid-like fashion. In Fig. 6 we plot a heatmap example for GEM trained on CIFAR-100-2, employing a buffer size of 500 ($2\%$ of the dataset size). We focus on small buffers, as the original works often mainly report results for such buffers. Our findings mirror our previous findings on the experience replay case: a large set of hyperparameters improves the algorithm’s performance, while prioritizing the removal of slowly learned examples proves advantageous given the buffer’s small size. Comparable outcomes across different algorithms, datasets, and class incremental learning scenarios are detailed in Apps. A,E.
Next, we compare continual learning methods with Goldilocks on popular continual learning benchmarks, including CIFAR-10-5, CIFAR-100-5, and TinyImageNet-10 datasets. The $quick$ and $slow$ hyperparameters are determined through RotNet training, as outlined in §3.3. Although these parameters might vary per task when $T>2$, we maintain consistent values across tasks to mitigate an excessive number of hyperparameters. Goldilocks consistently improves various continual learning algorithms, indicating its compatibility with diverse methods. These results are depicted in Fig. 7.
3.5 Comparing Goldilocks to other replay buffer sampling methods
While numerous replay buffer sampling methods have been proposed over the years, uniform sampling remains prevalent in practical implementation within continual learning algorithms. This preference may stem from the specific settings or algorithmic modifications required for other sampling methods to work well. These methods do not consistently improve different continual learning algorithms across datasets and buffer sizes. In contrast, Goldilocks sampling improves a wide range of continual learning algorithms across various buffer sizes and datasets.
We explored various replay buffer sampling strategies, including uniform sampling, max entropy, IPM [56], GSS [1], and Herding [44]. We trained $10$ networks on CIFAR-10-2 and CIFAR-100-2 datasets with buffer sizes of $1000$ and $10000$ (equivalent to $4\%$ and $40\%$ of the dataset size, respectively) using experience replay. Table 3.5 presents the mean test accuracy across all tasks for each case.
Goldilocks consistently outperforms all other sampling strategies across all scenarios, surpassing them in all four cases. In contrast, other sampling methods exhibit efficacy only in specific scenarios. For instance, Herding, which prioritizes examples with mean features close to class mean features, performs better than uniform sampling with smaller buffer sizes but worse with larger buffers. This might be because this preference for typical examples, which are learned early on, is more beneficial with smaller buffers. Conversely, GSS, which selects examples based on gradient distances, outperforms uniform sampling in scenarios with large buffers but performs suboptimally with smaller budgets. This bias towards selecting examples with large gradients favors examples learned slowly, which is more beneficial with large buffers.
(a) CIFAR-100-5
(b) TinyImagenet-10
Figure 7: Goldilocks vs. uniform sampling with various continual learning methods. Each bar shows the mean final test accuracy across all tasks for each method, with a uniform sampled buffer (traditionally done) in blue, and Goldilocks in orange. Error bars denote the standard error. (a) CIFAR-100-5 with a buffer size of $5\%$. (b) TinyImagenet-10 with a buffer of size $10\%$. Goldilocks significantly improves performance across all methods and datasets.
Table 1: Comparing replay buffer sampling methods. Each entry is the mean test accuracy across all tasks of 10 networks trained using experience replay on CIFAR-100-2. Bold denotes the method with the highest accuracy.
CIFAR-10-2 CIFAR-100-2
Buffer size $4\%$ $40\%$ $4\%$ $40\%$
Random 87.67 95.03 69.03 79.24
Max entropy 84.44 94.63 64.27 77.91
IPM 85.9 94.76 64.16 75.73
GSS 85.28 95.55 64.15 80.11
Herding 88.41 94.53 71.07 78.97
Goldilocks 88.99 96.03 71.53 80.62
4 Conclusion
Limitations. Goldilocks necessitates maintaining an epoch-wise classification matrix to compute the learning speed of each example. This estimation’s accuracy increases with the number of epochs completed. Consequently, in stream-based approaches, where each data point is encountered only once, Goldilocks cannot be applied. In such cases, alternative methods for approximating example complexity may be viable but fall beyond the scope of this study, leaving them for future work.
Another limitation arises in the potential lengthy computation time when fine-tuning hyperparameters in multi-task settings, as the $quick$ and $slow$ hyperparameters may differ between each subsequent task. However, in practice, these hyperparameters can be selected using simple heuristics and kept constant across tasks, as demonstrated in our training on multiple tasks (see Fig. 7).
Discussion. We show a correlation between the speed at which examples are learned and the likelihood of models forgetting these examples when learning new data. This correlation aligns closely with the simplicity bias, suggesting that models tend to forget more complex examples. Although measuring learning speed is a simple proxy for example complexity, its simplicity underscores the potential for developing more sophisticated methods to mitigate catastrophic forgetting, offering avenues for future research.
While most replay-based continual learning works focus on better utilization of the replay buffer, our study demonstrates the potential for buffer improvement, extending the scope of continual learning research beyond buffer utilization alone. By showing that Goldilocks can be implemented with minimal computational overhead, we highlight its applicability to future advancements in continual learning. Furthermore, this work underscores the mutual benefits of concurrent progress in better sampling of the replay buffer and improvements in buffer utilization, suggesting promising directions for future research.
Appendix
Appendix A Class incremental learning
In continual learning, a distinction is made between class incremental learning and task incremental learning. The key difference lies in test time: in task incremental learning, the model knows the task identity of the given examples, allowing it to use a specialized classifier that differentiates between classes within that task. Conversely, in class incremental learning, the model lacks task identity information at test time and must identify both the task and the class of the example, typically resulting in lower performance. Due to space constraints, we report only the task-incremental results in the main paper, as it is the more common benchmark. However, we also tested for class incremental settings in all our experiments. While the performance is lower, the qualitative outcomes remain consistent: buffer compositions that are good for task incremental learning are also good for class incremental settings and vice versa, and Goldilocks significantly enhances a variety of continual learning algorithms in both scenarios.
In Fig.8(a), we extend the results from Fig.4(b) to class incremental settings. We examine different buffer compositions by adjusting the $quick$ and $slow$ hyperparameters of Goldilocks and plot the mean final test accuracy across all tasks for 10 networks trained on CIFAR-100-2 with a $40\%$ buffer, using experience replay. The qualitative results are consistent: the same buffer compositions are effective for both learning scenarios. In Fig.8(b), we replicate Fig.6, applying the same experiment with GEM on CIFAR-100-2 using a $1\%$ buffer, confirming that the results align with those reported in the main paper.
Finally, in Fig.9, we reproduce the results from Fig.7, showing the performance of competitive continual learning algorithms when changing only the buffer sampling strategy to Goldilocks, across CIFAR-100-5, CIFAR-10-5, and TinyImageNet-10. In all datasets and with all continual learning algorithms, Goldilocks significantly improves the performance of the original methods by a significant margin.
(a) CIFAR-100-2, $40\%$ buffer
(b) GEM, CIFAR-100-2, $1\%$ buffer
Figure 8: Comparing different replay buffer compositions in the class incremental case. Each matrix compares the mean final test accuracy of 10 networks trained with different replay buffer compositions achieved by modifying Goldilocks’ $quick$ and $slow$ hyperparameters. The ($quick$, $slow$) = ($0\%$, $0\%$) case, marked by a black box, corresponds to uniform sampling. (a) Replicating Fig. 4(b) with class incremental learning (b) Replicating Fig. 6 with class incremental learning. In both cases, the results mirror the task-incremental case, although with lower performances.

### v2

@@ 2.4 Adding a Replay Buffer of Different Sizes
Figure 3: Impact of the buffer size on forgetting dynamics. (a-c) Similarly to Fig. 2(c), the first task’s binary matrices $M$, when adding replay buffer of varying buffer sizes: small ($1k$), medium ($3k$), and large ($12k$). The y-axes denote examples, and the x-axes denote epochs, indicating if the example has been classified correctly by the model at that epoch. The order in which examples appear is sorted by the examples’ learning speed. In all cases, examples that were learned faster are more likely to be remembered. With bigger buffers, the networks can remember gradually slower-to-learn examples. (d) The mean learning speed of remembered examples of models with different buffer sizes. Models with bigger replay buffers remember slower-to-learn examples.
@@
To evaluate replay-based methods, we incorporate a replay buffer in our experiments. Similarly to §2, we trained $10$ networks on CIFAR-100-2, with replay buffers of varying sizes ($1k$, $3k$, $10k$). In Figs. 3(a-c), we plot the epoch-wise classification matrix of the test data for each case. The examples appear in each matrix in the order of their learning speed, from slow (bottom) to quick (top). As expected, integrating a replay buffer mitigates some catastrophic forgetting, boosting the model performance on the first task, with larger buffers yielding better results. Nevertheless, similar to the scenario without a replay buffer $B=\emptyset$, a relationship exists between learning speed and catastrophic forgetting: networks tend to remember fast-learned examples while forgetting those learned later even when using a replay buffer. Notably, models with larger buffer sizes remember most of the examples that models with smaller buffer sizes do, while additionally remembering gradually slower-learned examples. Similar results occur when changing the network architecture, and when decreasing the dataset complexity (see App. C, D).
@@
We quantify the relationship between the buffer size and remembered examples’ learning speed. Training $10$ networks on CIFAR-100-2 with various buffer sizes, we plot the mean learning speed of the remembered examples for each buffer size. A strong correlation emerges ($r=-0.966$, $p\leq 10^{-6}$): models with smaller buffer sizes remember quickly-learned examples, whereas larger models with larger buffers enable the model also to remember slower-learned. These results are plotted in Fig. 3(d). Similar results were achieved across diverse datasets, architectures, hyperparameters, and continual learning algorithms, see App. C, D.
2.5 Different Buffer Compositions and Sizes
Replay-based continual learning methods often sample the replay buffer uniformly, treating quickly and slowly learned examples equally. Here, we analyze different replay buffer compositions, focused on examples learned at specific speeds. To focus the buffer on examples learned at a certain speed, we set 2 threshold, $quick$ and $slow$, and sample the buffer uniformly from all the examples that were learned slower than the $quick$ threshold and faster and the $slow$ threshold. This allows us to bias the replay buffer toward examples learned at certain speeds.
We compare replay buffer compositions for CIFAR-100-2 across different buffer sizes. Figs.4(a-b) show the mean final accuracy difference relative to uniform sampling, averaged over 10 runs. Compositions outperforming uniform sampling are in red, while those reducing performance are in blue. The black-boxed point $(quick,slow)=(0,0)$ represents the uniform baseline. Standard errors are plotted separately (App.E). In both cases, a wide range of buffer compositions (marked in red) significantly improves the final accuracy, suggesting that uniform sampling of the replay buffer is suboptimal.
While many buffer compositions are beneficial for both cases, we see that when the buffer size is smaller, it is more beneficial to remove slowly-learned examples, focusing mainly on the easier examples learned quickly by the models. In contrast, with larger buffers, the optimal focus in the buffer shifts towards harder, slowly-learned examples, where quickly-learned examples are removed.
@@ 2.5 Different Buffer Compositions and Sizes
In Fig. 4(c), we replicate this experiment on TinyImageNet-2, with a large buffer of size $20k$. Similar to CIFAR-100-2, focusing on examples learned midway through learning yields the best performances. Notably, across all cases, the range of thresholds that enhance performance is broad and continuous, indicating that improvements are not limited to specific hyperparameter sets. Similar results are achieved in the multi-task case, see App. C.5.
To assess the robustness of these results, we repeat the experiment from Fig. 4 under various learning settings, including changes to optimizers, architectures, learning rates, training durations, fine-tuning parameters, regularization techniques, and data splits. Across all cases, we observe consistent qualitative results. These experiments can be found in App. D.
(a) CIFAR-100-2, small buffer
(b) CIFAR-100-2, big buffer
(c) TinyImageNet-2, big buffer
Figure 4: Comparison of buffer compositions for various buffer sizes, removing slowly and quickly learned examples (see §3.1). The mean final accuracy difference between each composition and a randomly sampled buffer is shown, averaged over 10 runs. Red shades indicate improved performance over uniform sampling, while blue indicates worse performance. The baseline (uniformly sampled buffer) appears at ($0\%$, $0\%$), marked with a black box. (a) CIFAR-100-2 with a small $1k$ buffer: removing slower-to-learn examples is more beneficial. (b) CIFAR-100-2 with a large $10k$ buffer: removing faster-to-learn examples is more beneficial. (c) TinyImageNet-2 with a $20k$ buffer. Across all scenarios, diverse buffer compositions are beneficial, with smaller buffers benefiting more from faster-to-learn examples.
2.6 How Subsequent Tasks Affect Buffer Composition
We empirically examine how subsequent tasks impact the ideal composition of a replay buffer for a given task. We find that regardless of the similarity or dissimilarity between subsequent tasks and the original task, the optimal replay buffer composition remains largely independent and consistent. We conduct experiments with three distinct tasks denoted A, B, and C. By training the model on task A and subsequently introducing either task B or task C while keeping the original task unchanged, we analyze the ideal replay buffer composition under varying subsequent tasks.
@@ 2.6 How Subsequent Tasks Affect Buffer Composition
We initially focus on tasks from the same datasets. We pick task A to be the first 25 classes of CIFAR-100, task B to be classes 26 to 50, and task C to be classes 51 to 75. In Fig. 5, we train various replay buffer compositions when either task B or task C is after task A. We plot the difference between the mean final accuracy of each buffer composition to a uniformly sampled buffer. Each composition was repeated 10 times. The consistent qualitative results indicate that good buffer compositions for one case are also good for the other, and vice versa. These results suggest that the ideal replay buffer composition of task A is robust to differences between tasks B and C in this context.
@@
Further, we explore scenarios involving either changing the classification task or replacing it with memorization by training on random labels. We adopt the RotNet approach (Gidaris et al., 2018) to change the classification task. We pick tasks A and B to be different subsets of CIFAR-100, while task C involves a rotation classification task, where each example in task A is randomly rotated between $90\degree,180\degree$, or $270\degree$ degrees, with the task being to classify these rotations regardless of the original label. For the random case, tasks A and B remain unchanged, while task C involves the same subset of task B, but with labels chosen uniformly at random, as suggested in Zhang et al. (2021). In both scenarios, we observe consistent results: the ideal replay buffer composition of task A remains robust to differences between tasks B and C in both contexts. Further details and heatmaps for the results are available in App. F.
(a) (A)$\rightarrow$(B)
(b) (A)$\rightarrow$(C)
Figure 5: Buffer compositions when training on the same initial task, followed by different subsequent tasks. We use three subsets of CIFAR-100: A, B, and C (see §2.6). (a) Task A followed by B. (b) A followed by C. Despite the differences in subsequent tasks, the buffer compositions for the original task similarly mitigate catastrophic forgetting.
3 Speed-Based Sampling (SBS)
Above, we observed a distinctive behavior of neural networks trained in a continual learning setting: when exposed to new data, networks tend to forget previously learned data, with examples that were learned more quickly being the least likely to be forgotten. This section demonstrates that this behavior has practical implications and can be leveraged to mitigate catastrophic forgetting in CL.
We propose Speed-Based Sampling (SBS), a novel sampling strategy for replay buffers that replaces traditional random sampling which is common in various CL methods. SBS prioritizes examples with specific learning speeds, allowing the replay buffer to account for the susceptibility of examples to forgetting. This allows the underlying CL algorithm to focus on the most relevant examples, effectively reducing catastrophic forgetting.
3.1 Method Definition
Speed-Based Sampling (SBS) is a sampling strategy for the replay buffers in buffer-based CL at the end of each task. It is compatible with any buffer-based CL algorithm.
During training on task $t$, SBS calculates the classification matrix (defined in §2.1) of task $t$, computing at the end of the task the learning speed of each example in $\mathcal{D}_{t}$. Once training on task $t$ is completed, SBS creates a filtered dataset $\mathcal{D}^{\prime}_{t}$ by removing the $q$ quickest-learned examples and $s$ slowest-learned examples from $\mathcal{D}_{t}$. Examples for the replay buffer are then sampled uniformly from $\mathcal{D}^{\prime}_{t}$.
The hyperparameters $q$ and $s$ can be adjusted based on the user’s requirements (see below). In scenarios where the buffer size is limited and examples need to be removed (e.g., when handling multiple tasks), examples from previous tasks can be removed from the buffer at random.
Pseudocode for integrating SBS with any buffer-based CL method is provided in Algorithm 1.
Algorithm 1 Training $CL$ method with SBS.
Input: $\mathcal{D}_{t}$, $|B|$, $E$, amount of quick/slow to remove $q,s$. Output: buffer of size $|B|$
  Classification_Matrix $\leftarrow$ ${0}^{E\times|\mathcal{D}_{t}|}$
  for $e=1,...,E$ do
     Train the model $f$ one epoch using the $CL$ method
     for $(x_{i},y_{i})\in\mathcal{D}_{t}$ do
        Classification_Matrix$[e,i]\leftarrow\mathds{1}[f(x_{i})=y_{i}]$
     end for
  end for
  $Speed$leftarrow Mean(Classification_Matrix, axis=0)
  $\mathcal{D}_{t}^{{}^{\prime}}\leftarrow$ remove $q\%$ quickest and $s\%$ slowest from $\mathcal{D}_{t}$
  return $|B|$ examples sampled uniformly from $\mathcal{D}_{t}^{{}^{\prime}}$
Choosing $q$ and $s$. The choice of $q$ and $s$ is crucial for the performance of SBS. In our experiments, we determined these hyperparameters by training a RotNet auxiliary task on the same dataset, without using additional data points. Specifically, after training the CL method on $\mathcal{D}_{t}$, we created a rotated version of the dataset by randomly rotating each image by $0\degree$, $90\degree$, $180\degree$, or $270\degree$. Then we trained the network to classify the rotation of each example. A coarse grid search was performed over $q$ and $s$ values, selecting the ones that yielded the best performance on this task.
While effective in practice, this approach requires training the model multiple times, which can be computationally expensive for larger datasets. However, the smoothness of the parameter space (see Fig.4) enables a coarse grid search to be sufficient. Moreover, since the auxiliary task is relatively simple, each training run is significantly faster than training the model from scratch on the full dataset $\mathcal{D}_{t}$, as fewer epochs are needed. In addition, the method is robust to hyperparameter choices: a broad range of values for $q$ and $s$ tend to yield improvements, making heuristic settings often adequate. For example, setting $q=s=20\%$ consistently enhances performance across all evaluated datasets and generalizes well to unseen ones. This is evident in Figs. 4(a–c), 9(a–b), 16(a–b), 17(a–b), 18(a–c), 19(a–c), 20(a–c), 21(a–c), and 24(a–c), where the point corresponding to $q=s=20\%$ consistently significantly outperforms random sampling, even if it does not always achieve the optimal result. Table 3.2 further compares the performance of SBS with and without hyperparameter tuning. Even in the fixed setting (SBS-fix, with $q=s=20\%$), the method consistently improves performance across all datasets and buffer sizes tested.
Further refinements can also be guided by task-specific rules. Based on the results of §2.5, if the task is relatively easy for the model (resulting in less catastrophic forgetting), it is beneficial to focus more on slower-to-learn examples, increasing $q$ and decreasing $s$. Conversely, for more challenging tasks where forgetting is pronounced, decreasing $q$ and increasing $s$ is more appropriate.
Running time. Given a set of $q$ and $s$ parameters, SBS introduces minimal computational overhead to the underlying CL method. The classification matrix required by SBS is computed using the classifications of each example during each epoch, which are already part of the forward pass. While the matrix has a size of $E\times|\mathcal{D}_{t}|$, potentially large for many epochs, this memory cost can be reduced by maintaining only a running mean for each example, lowering the storage requirement to $|\mathcal{D}_{t}|$. Additionally, as SBS samples examples randomly from a smaller data pool, the sampling process has a similar runtime to standard random sampling.
3.2 SBS Improves Various CL Algorithms